# IG Parser: A Software Package for the Encoding of Institutional Statements using the Institutional Grammar

### v1

@@ Abstract
This article provides an overview of IG Parser, a software that facilitates qualitative content analysis of formal (e.g., legal) rules or informal (e.g., socio-normative) norms, and strategies (such as conventions) – referred to as institutions – that govern social systems and operate configurally to describe institutional systems. To this end, the IG Parser employs a distinctive syntax that ensures rigorous encoding of natural language, while automating the transformation into various formats that support the downstream analysis using diverse analytical techniques. The conceptual core of the IG Parser is an associated syntax, IG Script, that operationalizes the conceptual foundations of the Institutional Grammar, and more specifically Institutional Grammar 2.0, an analytical paradigm for institutional analysis. This article presents the IG Parser, including its conceptual foundations, syntactic specification of IG Script, alongside architectural principles. This introduction is augmented with selective illustrative examples that highlight the use and benefit associated with the tool.
@@ B.1 Institutions
Understanding the functioning of social systems, whether at group, organizational, societal, or international level, relies on the ability to extract and analyze the rules that coordinate behavior, mitigate (or sometimes provoke) conflict and drive the necessary cohesion in diverse and open societies. Those institutions can take various forms, manifesting as observed behavioral patterns of collective action, and of course in written form, be it in the form of instructions, regulations, laws, or other form of private or public policy. One way of making such – typically qualitative – information analytically accessible is to employ content analysis techniques. While text analysis is well established for capturing features of natural language (e.g., based on dependency tree parsing in the context of Natural Language Processing), language used for institutions is distinct, since a) not all language expresses rules, and b) institutional language displays distinctive functional properties that can be expressed in a uniform syntax whose features correspond to distinctive semantic features of rules. One approach to capture the essential features of such language at a fine-grained level is the Institutional Grammar [1, 2], a paradigm for institutional analysis that has found broad and long-standing disciplinary application, covering areas such as legal analysis [7], political science, specifically public policy and administration [8, 9, 10, 11, 6], socio-ecological systems analysis [12], as well as computational social scientists interested in the study of institutions [13, 14, 15]. [^14] For comprehensive reviews of the application landscape, please consult [16] and [17].

### v2

@@ Abstract
This article provides an overview of IG Parser, a software that facilitates qualitative content analysis of formal (e.g., legal) rules or informal (e.g., social) norms, and strategies (such as conventions) – referred to as institutions – that govern social systems and operate configurally to describe institutional systems. To this end, the IG Parser employs a distinctive syntax that ensures rigorous encoding of natural language, while automating the transformation into various formats that support the downstream analysis using diverse analytical techniques. The conceptual core of the IG Parser is an associated syntax, IG Script, that operationalizes the conceptual foundations of the Institutional Grammar, and more specifically Institutional Grammar 2.0, an analytical paradigm for institutional analysis. This article presents the IG Parser, including its conceptual foundations, the syntax specification of IG Script, and its architectural principles. This overview is augmented with selective illustrative examples that highlight its use and the associated benefits.
@@ B.1 Institutions
Understanding the functioning of social systems, whether at group, organizational, societal, or international level, relies on the ability to extract and analyze the rules that coordinate behavior, mitigate (or sometimes provoke) conflict and drive the necessary cohesion in diverse and open societies. Those institutions can take various forms, manifesting as observed behavioral patterns of collective action, and of course in written form, be it in the form of instructions, regulations, laws, or other form of private or public policy. One way of making such – typically qualitative – information analytically accessible is to employ content analysis techniques. While text analysis is well established for capturing features of natural language (e.g., based on dependency tree parsing in the context of Natural Language Processing), language used for institutions is distinct, since a) not all language expresses rules, and b) institutional language displays distinctive functional properties that can be expressed in a uniform syntax whose features correspond to distinctive semantic features of rules. One approach to capture the essential features of such language at a fine-grained level is the Institutional Grammar [1, 2], a paradigm for institutional analysis that has found broad and long-standing disciplinary application, covering areas such as legal analysis [7], political science, specifically public policy and administration [8, 9, 10, 11, Kianpour2024AnalysisGood], socio-ecological systems analysis [12], as well as computational social scientists interested in the study of institutions [13, 14, 15]. [^14] For comprehensive reviews of the application landscape, please consult [16] and [17].